OpenStudy (anonymous):

evaluate the following expression (4z)^2-44 when z=3

OpenStudy (anonymous):

100

OpenStudy (anonymous):

just plug3 in for z, = 100

OpenStudy (anonymous):

=(4*3)^2-44 =144-44=100

OpenStudy (anonymous):

144-44=100

OpenStudy (anonymous):

(4x3)^2 -44 12^2 -44 144-44 =100